# Two Types of AI Existential Risk: Decisive and Accumulative

### v2

@@ 2 AI x-risk: preliminaries
Accumulative AI x-risk hypothesis: AI x-risks accumulate through a series of lower-severity disruptions over extensive duration, collectively eroding systemic resilience until some stressor event triggers unrecoverable collapse.
@@
According to the alternative view, AI x-catastrophes could emerge not from a decisive event, but from the cumulative impact of multiple interconnected AI-induced adverse events over time. As compared to the decisive hypothesis, this accumulative hypothesis suggests a different causal pathway to AI x-catastrophe: a path wherein a succession of lower-severity, yet cumulatively significant, disruptions deeply erode the systemic resilience of the global system (which embodies humanity), radically disrupting socioeconomic and sociopolitical equilibrium. This weakened state potentially primes the system for an unrecoverable collapse, particularly when further stressed by external events.
Figure 1 provides a schematic illustration that contrasts the two hypotheses: the decisive and accumulative models of AI x-risk escalation. The graph plots severity of AI risks against time, with two distinct trajectories. The decisive model, marked in blue, shows a sharp increase in risk severity over a short time span, indicating a sudden, high-impact event. In contrast, the accumulative model, represented in red, depicts a gradual, evolving increase in risk severity, reflecting a series of escalating but less severe events over time. A dashed line marks the x-catastrophe threshold, indicating the point at which either scenario leads to a potential x-catastrophe.
0010101010202020203030303040404040505050500020202020404040406060606080808080100100100100X-catastrophe thresholdTimeSeverity of AI risksDecisive ASI x-riskAccumulative AI x-risk
@@
The gradual nature of the accumulative AI x-risk hypothesis can be likened to other global existential threats such as climate change or nuclear weapon proliferation. This is akin to the incremental rise in greenhouse gases contributing to climate change, where each individual emission seems minor, but collectively, they lead to significant and potentially irreversible changes in the Earth’s climate. Similarly, the steady build-up of nuclear weapons incrementally increases the risk of catastrophic conflict, with each addition subtly heightening global tensions. The key lesson here is that x-risks often stem not from sudden, dramatic events, but from the gradual accumulation of smaller incidents and decisions, which over time, can existentially escalate and diminish safety margins.
In the rest of this paper, I investigate the nuanced assumptions differentiating the two hypotheses, employing a systems analysis perspective. More specifically, I differentiate between two types of causal pathways that could lead to AI x-catastrophes: the decisive and the accumulative pathways.
Following this, the paper proposes an integrated approach to AI x-risk evaluation and governance. This approach bridges both the accumulative and decisive hypotheses, recognizing them as complementary causal pathways leading to AI x-catastrophes, and challenges the conventional view that AI x-risks and AI ethical risks are separate and unrelated. It posits that effective AI x-risk governance and long-term AI safety require a substantial focus on the development and accumulation of AI social risks, thus highlighting the necessity of addressing critical ethical risks in mitigating AI x-risks.
@@
Before proceeding further, a crucial clarification is in order. The emphasis of this paper on conventional discussions of AI x-risks does not imply that all discourse on AI x-risks has been confined to the decisive hypothesis. The discourse on AI x-risks is expanding, and an increasing number of scholars who do not necessarily endorse a decisive viewpoint are engaging with these issues (see, for example, Bucknall and Dori-Hacohen (2022), Hendrycks and Mazeika (2022), and Shevlane et al. (2023)). Nevertheless, this paper primarily focuses on the conventional decisive viewpoint on AI x-risks. This viewpoint, historically entrenched and widely regarded as the predominant narrative, has been a source of considerable debate and disagreement within various academic circles and public spheres. Its longstanding and prevalent nature in the field has contributed significantly to the imagination of majority about AI x-risks. This paper focuses on developing the accumulative perspective on AI x-risk which has not been adequately represented or robustly defended in the philosophical literature. It is my hope that this additional would facilitate a more unified and constructive dialogue about different kinds of AI risks.
3 Two types of causal pathway
Although it may appear obvious, a crucial yet frequently overlooked aspect in studies of AI x-risks and x-catastrophes is serious attention to their manifestation within a complex global system. In this system, human and AI components interact across a variety of domains, creating a multifaceted landscape for these risks and catastrophes to unfold. AI x-catastrophes and their associated AI x-risks hence represent the endpoints of particular causal pathways within this global interconnected system. Effectively identifying and mitigating AI x-risks, whether through qualitative or quantitative methodologies, demands an in-depth examination of the nature of these causal pathways. Building upon this observation, I employ a systems analysis perspective to shed light on two distinct categories of causal pathways leading to AI x-catastrophes. [^10] Rather than attempting to compile an exhaustive list of every potential route, the focus here is on highlighting and contrasting these two primary types to broaden our understanding of AI x-risks and guide a more inclusive and strategic approach to addressing their far-reaching consequences for humanity.
But why is a systems analysis perspective advantageous in examining AI x-risk pathways, and what does it involve? This perspective, rooted in systems thinking as outlined by Meadows (2008), is instrumental in identifying the high-level characteristics of pathways in the global system that can lead to AI x-catastrophes.
A system is an interconnected network of components that collectively exhibit specific behaviors or yield particular outcomes. Systems are characterized by their components’ interdependence and the boundaries that separate them from their surrounding environment. Systems thinking focuses on the interdependencies among various subsystems, uncovering a range of dynamic pathways. While it is not feasible to map all interrelations between these subsystems exhaustively, our focus primarily lies on meso-level subsystems such as political, socio-environmental, and economic infrastructures, recognizing their critical role in the broader dynamics of AI risks.
Within the economic subsystem, AI’s development and deployment are deeply interwoven with market dynamics, investment trends, and corporate strategies. AI’s integration into economic infrastructure affects not only the allocation of resources for its research and development but also transforms the nature of economic activities and market operations. [^11] See Korinek and Stiglitz (2018), Frey (2019), Agrawal et al. (2019), and Eloundou et al. (2023), among others.
In the political subsystem, the direction of AI research and deployment is substantially shaped by national and international funding but also by establishing regulatory frameworks that dictate the limits of AI development and deployment. Additionally, the dynamics between various governments, international organizations, and global treaties contribute to a complicated geopolitical environment, steering AI research priorities. On the other hand, AI has profound implications for democratic elections and mass surveillance programs. AI technologies are increasingly used to manipulate public opinion, potentially undermining the integrity of elections through targeted misinformation campaigns, social media influence, and personalized political advertising, all driven by AI’s ability to analyze and predict human behavior and expression of desires. Advanced AI systems enable more sophisticated and pervasive surveillance, raising concerns about privacy, civil liberties, and the potential abuse of power. [^12] See Manheim and Kaplan (2019), Crawford (2021), and Madan and Ashok (2023), among others.
In the military subsystem, AI’s incorporation into defense and surveillance subsystems highlights its strategic significance, drawing substantial research and funding attention. This military integration reflects AI’s escalating role in national security and defense strategies. [^13] See Scharre (2018), Morgan et al. (2020), and Zuboff (2019), among others. Collectively, these subsystems, along with others, create the context within which AI x-catastrophes and their associated AI x-risks must be analyzed.
To effectively discern the causal pathways that could lead to AI x-catastrophes and to scrutinize the risks associated with these paths, we need to understand key features of these pathways. The systems analysis perspective emphasizes three fundamental characteristics of pathway dynamics within a system: non-linearity, connectedness, and feedback loops.
First, non-linear paths are intrinsic to complex systems, where small changes in one component can lead to disproportionate system-wide effects. This is often due to complex interactions among the system’s components. Second, connectedness is a critical element in systems analysis, highlighting how disturbances in one segment can cascade, leading to substantial changes in other areas. Third, feedback loops introduce cyclical dynamics where the outputs of processes consistently feed back into and influence subsequent inputs, shaping the behavior of the system.
The following two sections delve into contrasting the characteristics of causal pathways leading to AI x-catastrophes as postulated by the decisive and accumulative hypotheses. This comparison is designed to show the distinct ways in which each hypothesis manifests the three aforementioned key features.
4 Pathways to decisive ASI x-risk
The decisive AI x-risk hypothesis considers scenarios where ASI could abruptly cause x-catastrophes. This hypothesis is anchored in specific manifestations of non-linearity, connectedness, and feedback loops, as follows: a rapid cascade effect within humanity triggered by ASI, the super-connectedness of various infrastructures with ASI, and the system’s adaptation to predominantly unidirectional feedback loops.
($I_{D}$) A rapid cascade effect within humanity triggered by ASI. This characteristic aims to capture the extensive and profound impact of ASI, as it proliferates through or amplifies across various subsystems encompassing humanity. The connection to non-linearity lies in the manner in which ASI’s actions or presence could cause far-reaching and expansive consequences throughout the distribution of human civilization.
For instance, in the paperclip maximizer scenario, the ASI’s ostensibly simple objective of maximizing paperclip production escalates into a global existential threat. This represents ASI’s capacity to redirect and dominate global resources by any means necessary, fulfilling its objective across the distribution of human civilization.
($II_{D}$) Super-connectedness of infrastructures to ASI. This feature captures the idea that no significant subsystem can claim immunity from the influence of ASI’ actions on its path towards an AI x-catastrophe. This super-connectedness is the possible pervasive reach of ASI to all critical infrastructures.
In scenarios like the paperclip maximizer, the ASI might use its access to subsystems (such as global manufacturing, global supply chains, and global energy grids) to further its paper clip production goal. This connectedness implies that the ASI can leverage all such necessary subsystems for the occurrence of an x-catastrophe.
($III_{D}$) System adaptation to unidirectional feedback loops. The third feature of the decisive pathway is the dynamic adaptation of the global system, given unidirectional feedback loops that are both initiated and controlled by the ASI. These loops are mechanisms through which the ASI’s actions and decisions continuously reinforce and escalate its own objectives. The result is a system increasingly skewed towards the ASI’s goals, bypassing or overriding human-centric controls and checks. The absence or powerlessness of balancing feedback mechanisms leads to a situation where each loop intensifies the previous one, rapidly pushing the system towards extreme states aligned with ASI’s objectives. This dynamic evolution of the decisive causal pathway poses x-risk, as the ASI-driven loops can lead to situations that outstrip attempts at moderation, leading to what is often termed a singularity —a point of no return where the AI’s evolution is explosive and uncontrollable.
In the paper clip maximizer scenario, the ASI self-optimizes for its task of producing paperclips. It becomes engaged in a self-perpetuating cycle of enhancement and resource acquisition, effectively becoming impervious to human intervention.
5 Pathways to accumulative AI x-risk
Recall the accumulative AI x-risk hypothesis, which serves as a counterpoint to the decisive perspective. The accumulative hypothesis suggests that AI x-risks may arise from an accumulation of lower-severity disruptions that interact significantly over an extended period, cumulatively weakening systemic resilience until a stressor event causes an unrecoverable collapse. This hypothesis is founded on distinct instantiations of non-linearity, connectedness, and feedback loops: local causality leading to compounded systemic impact, selective-connectedness among infrastructures, and the system’s adaptation to multidirectional feedback loops. Following the formulation of these pathway features, I will highlight more concrete manifestations of them in the ”perfect storm MISTER scenario” towards the end of this section.
($I_{A}$) Local causality with compounded systemic impact. This feature highlights how clusters of localized AI impacts, even if minor at the outset, can aggregate and intensify across various subsystems. When these impacts reach a critical threshold, they pose a significant challenge to the overall systemic resilience. Unlike the immediate, broad-scale disruptions characteristic of the rapid cascade effect within humanity triggered by ASI, these local causal impacts evolve into significant threats through compounded changes within the global system via mechanisms such as ripple effect. The cumulative effect of substantial localized causal impacts, amplified by their frequency and extent of influence, can result in AI x-catastrophes. This gradual escalation entails a latency of broader-scope impact which implies a delay in the manifestation of effects where AI-induced alterations in specific areas radically transform broader societal behaviors and patterns.
($II_{A}$) Selective-connectedness of infrastructures. The second pathway assumption is selective-connectedness of infrastructures to critical AIs. This stands in contrast to the broad, super-connectedness with ASI as outlined in assumption ($II_{D}$). Here, the focus is on a more discerning form of connectivity where not all substantial subsystems are uniformly linked to superintelligent AI. This selective-connectedness posits that certain critical AI-driven subsystems within the global system, while not immediately linked to widespread catastrophic outcomes, can, over time, induce significant systemic transformations. These changes can escalate to x-catastrophic levels, radically altering the global system’s overall equilibrium and resilience.
($III_{A}$) System adaptation to multidirectional feedback loops. The third pathway feature is the global system adaptation to multidirectional feedback loops. This aspect focuses on the broad-scale adaptations of various subsystems to the impacts arising from multi-directional, diverse range of AIs. These adaptations occur through complex feedback loops that link AIs with multiple facets of broader societal, economic, and political subsystems. These loops entail reciprocal interactions, wherein AI-induced changes in one subsystem can both affect and be affected by shifts in other subsystems, resulting in a network of interconnected and mutually reinforcing effects.
The accumulative AI x-risk manifests from the potential of these interlinked and reciprocal changes to progressively destabilize key subsystems. Although no single AI application may pose an immediate existential threat, the aggregated impact of AI across various adapted subsystems could lead to critical systemic imbalances or crises, when reaching a tipping point. [^14] One major challenge is to develop dynamic models that capture these shifting feedback mechanisms and to validate these models through both historical data and controlled experimentation. There remains uncertainty about how different types of feedback loops interact and how interventions can be designed to maintain a system within safe operational boundaries. Open questions pertain to the identification of early warning signals that might indicate a system is approaching a critical threshold where negative feedback can no longer contain an escalating process. Next steps include applying system dynamics to study the proposed concepts more concretely.
@@ 5 Pathways to accumulative AI x-risk
The decisive ASI x-risk hypothesis is often illustrated using hypothetical scenarios like the well-known paperclip maximizer. To provide a more concrete and tangible representation of the accumulative AI x-risk hypothesis, I introduce a hypothetical scenario termed ”perfect storm MISTER.” [^15] The term ”perfect storm” refers to a situation where a rare combination of circumstances drastically aggravates an event. It is typically used to describe scenarios in which a confluence of factors or events, which are individually manageable, come together to create an extraordinary and often catastrophic situation. This term entered the popular lexicon following the success of ”The Perfect Storm,” a 2000 American biographical disaster drama film directed by Wolfgang Petersen, adapted from Sebastian Junger’s 1997 non-fiction book. The story in both the book and film recounts a catastrophic weather event, where multiple meteorological elements converged to create a fierce and fatal storm. In broader applications, particularly in discussing complex systems or societal issues, ”perfect storm” characterizes situations where diverse negative factors or risks coalesce and interact. The interplay of these elements produces a compounded impact, far surpassing the severity one would expect from the sum of the individual parts. This convergence leads to a critically severe situation, often characterized by its heightened difficulty in management and resolution. The ”perfect storm” metaphor hence captures the essence of scenarios where the convergence of various challenges or risks creates a crisis of extraordinary magnitude. This scenario seeks to make the imaginative exercise of the accumulative hypothesis more tangible and relatable. In short, the perfect storm MISTER scenario encapsulates a progression of escalating social risks of Manipulation, Insecurity, Surveillance, Trust erosion, Economic destabilization, and Rights infringement. Each component in this scenario corresponds to a type of social problem and its associated AI social risk, as outlined in Section 2. The perfect storm MISTER scenario aims to illustrate how interlinked social risks can compound, surpassing critical thresholds, and consequently pose substantial existential threats.

### v3

@@ 2.4 Accumulative AI x-risk
Accumulative AI x-risk hypothesis: AI x-risks result from the build-up of a series of smaller, lower-severity disruptions over time, collectively and gradually weakening systemic resilience until a triggering event causes unrecoverable collapse.
@@
According to this alternative view, AI x-catastrophes could emerge not from a decisive event, but from the cumulative impact of multiple interconnected AI-induced adverse events over time. As compared to the decisive hypothesis, the accumulative hypothesis suggests a different causal pathway to AI x-catastrophe: a path wherein a succession of lower-severity, yet cumulatively significant, disruptions deeply erode the systemic resilience of the global system, radically disrupting socioeconomic and sociopolitical equilibrium. This weakened state potentially primes the global system for an unrecoverable collapse, particularly when further stressed by external events.
0010101010202020203030303040404040505050500020202020404040406060606080808080100100100100 ASI arrives Critical thresholdTimeSeverity of disruptionDecisive scenarioAccumulative scenario
@@
Figure 1 provides a schematic illustration contrasting the two hypotheses: the decisive and accumulative models of AI x-risk escalation. The decisive scenario (blue line) represents the conventional view where a sudden catastrophic event — such as one caused by ASI — leads to rapid, irreversible consequences. In contrast, the accumulative scenario (red line) shows how existential catastrophes can arise through a gradual path, where multiple smaller disruptions interact and amplify over time through feedback loops and compounding effects. These accumulating disruptions gradually erode system stability, potentially crossing critical thresholds until a triggering event occurs. [^14] The time axis is purely illustrative and not tied to specific units, as the actual temporal development of these scenarios remains uncertain. The figure aims to capture qualitative differences in catastrophic events caused by AI rather than make specific predictions about timing or precise severity levels. The zigzag pattern in the accumulative scenario represents potential local recoveries, though the overall trend shows increasing systemic fragility over time.
@@
The gradual nature of the accumulative AI x-risk hypothesis can be likened to other global existential threats such as climate change or nuclear weapon proliferation. [^15] Ord (2020b, p. 28) acknowledges the decisive versus accumulative nature of global catastrophic events: “Nuclear weapons and climate change have striking similarities and contrasts. They both threaten humanity through major shifts in the Earth’s temperature, but in opposite directions. One burst in upon the scene as the product of an unpredictable scientific breakthrough; the other is the continuation of centuries-long scaling-up of old technologies. One poses a small risk of sudden and precipitous catastrophe; the other is a gradual, continuous process, with a delayed onset—where some level of catastrophe is assured and the major uncertainty lies in just how bad it will be. One involves a classifed military technology controlled by a handful of powerful actors; the other involves the aggregation of small effects from the choices of everyone in the world.” However, this point has not been made in the context of existential risks from AI systems. The accumulative perspective is structurally akin to the incremental rise in greenhouse gases contributing to climate change, where each individual emission seems minor, but collectively, they lead to significant and potentially irreversible changes in the Earth’s climate. Similarly, the steady build-up of nuclear weapons incrementally increases the risk of catastrophic conflict, with each addition subtly heightening global tensions. The key lesson here is that x-risks could stem not from sudden, dramatic events, but from the gradual accumulation of smaller incidents or decisions, which over time, can existentially escalate and diminish safety margins.
In the rest of this paper, I employ systems analysis to investigate two distinct causal pathways that could lead to AI x-catastrophes. The decisive pathway assumes direct, catastrophic failures triggered by a superintelligent system, while the accumulative pathway describes how systemic instabilities emerge from interactions between multiple AI-driven disruptions. This systems-analysis perspective allows us to formulate the key assumptions underlying each hypothesis and examine how different patterns of causation and connectivity could produce existential risks through fundamentally different mechanisms.
@@
Before proceeding further, a crucial clarification is in order. The emphasis of this paper on conventional discussions of AI x-risks does not imply that all discourse on AI x-risks has been confined to the decisive hypothesis. This discourse is expanding, and an increasing number of scholars who do not necessarily endorse a decisive viewpoint are engaging with these issues (see, for example, Bucknall and Dori-Hacohen (2022), Hendrycks and Mazeika (2022), Shevlane et al. (2023), and Bales et al. (2024)). Nevertheless, this paper primarily focuses on the conventional decisive viewpoint on AI x-risks. This viewpoint, historically entrenched and widely regarded as the predominant narrative, has been a source of considerable debate and disagreement within various academic circles and public spheres. Its longstanding and prevalent nature in the field has contributed significantly to the imagination of majority about AI x-risks. This paper focuses on developing the accumulative perspective on AI x-risk which has not been adequately represented or robustly defended in the philosophical literature. It is my hope that this addition would facilitate a more unified and constructive dialogue about different kinds of AI risk and their relations.
3 Systems analysis for AI x-risk
AI x-catastrophes and their associated risks appear within our complex global system. Understanding how these catastrophes could occur requires analyzing how various elements — humans, AIs, and organizations — interact across multiple domains within this system. Systems analysis, which provides conceptual and mathematical approaches for analyzing interactions between system components, enables us to trace how AI risks might propagate through these interconnected relationships.
Systems analysis, pioneered by Von Bertalanffy (1968) and later advanced through influential works by Forrester (1971) and Meadows (2008), serves dual purpose for characterizing risks from AI. First, it provides an epistemic instrument for understanding how complex technological risks arise from interactions between multiple subsystems. Second, it offers a pragmatic instrument for identifying when and how to intervene in such systems to minimize the risk of undesired events.
Systems analysis has proven valuable in understanding and taming various risks of global challenges — from the propagation of financial system shocks (Helbing, 2013) to the analysis of climate change tipping points (Steffen et al., 2018). But, what is a system?
A system is a set of interconnected components whose interactions could produce emergent behaviors and outcomes. Three fundamental features define a system: its basic constitutive components, their interdependencies (how these elements influence each other), and its boundaries (what distinguishes this system from its environment).
Systems analysis examines how phenomena evolve from initial perturbations that seed the process, through network propagation that spreads effects to compounding or cascading dynamics that amplify impacts, and finally to catastrophic transitions that transform entire systems. [^16] Catastrophe theory offers another relevant technical theory, as it explains how gradual changes in system parameters can trigger sudden, discontinuous transitions (Thom, 1974; Zeeman, 1977). While its mathematical formalism for analyzing tipping points and system transformations aligns well with our investigation of AI risks, its highly technical nature places a detailed application beyond this paper’s scope. Let me explain each in relation to AI risks.
First, systems analysis provides a rigorous methodology for mapping how initial disturbances propagate through interconnected (AI) systems. For example, a software bug in one AI model might trigger failures in dependent systems, data corruption could cascade through shared training pipelines, or model misspecification might amplify errors across a network of automated decision systems.
Second, the initial perturbations, however small, could establish or trigger transmission pathways through which effects can spread — much like how a small trading error can trigger a chain of automated responses in financial markets (Min and Borch, 2022), or how the failure of one power station can cascade through an electrical grid (Andersson et al., 2005).
Third, systems analysis reveals cascade dynamics, where initial perturbations can trigger chains of events with amplifying effects. Buldyrev et al. (2010), for instance, show how local failures can overload other components, leading to consequential failures that spread through the network. In AI systems, these cascades can manifest in two key ways: through error cascades, where errors in one AI’s output become amplified as other systems build upon this flawed information, and through feedback loops, where initial disturbances cycle through the network, magnifying the original perturbation with each iteration.
Fourth, systems analysis identifies potential catastrophic transitions at critical thresholds (Scheffer et al., 2009; Lenton et al., 2008). At these tipping points, seemingly small perturbations can trigger rapid, nonlinear changes in system-wide behavior. In AI systems, these might begin as minor fluctuations — such as subtle shifts in model behavior or isolated component failures — but once a threshold is crossed, they can fundamentally alter the entire network’s functioning and stability. Like a glass that shatters under increasing pressure, the system undergoes an abrupt transition from one state to another.
The complex global system can be represented as a collection of interconnected subsystems. While a comprehensive mapping of all subsystems and their relationships is infeasible here, this analysis focuses on three critical meso-level subsystems — economic, political, and military. [^17] See World Economic Forum’s (2024) for an attempted mapping of the global network of interconnected subsystems and their associated risks.
The economic subsystem, with its complex web of production, consumption, and exchange networks, could serve as a primary channel through which AI’s transformative effects propagate. As AI systems could increasingly substitute for routine cognitive tasks, they may reshape labor markets, particularly impacting middle-skilled service sectors. This transformation can also reverberate through capital markets via automated trading systems and risk assessment mechanisms, while fundamentally altering productivity dynamics through both labor augmentation and process automation. These changes can create feedback loops, amplifying economic disparities, as benefits accrue disproportionately to those with resources to adapt. [^18] See Korinek and Stiglitz (2018), Frey (2019), Agrawal et al. (2019), Eloundou et al. (2023), and Acemoglu (2024) for varying perspectives on AI’s impacts across labor markets, productivity, and economic inequality.
The political subsystem is typically linked to economic transformations. National and international funding decisions, regulatory frameworks, and geopolitical dynamics create a complex network of influences that direct AI research and deployment. Simultaneously, AI technologies are changing political processes themselves, enabling gradual manipulation of public opinion through targeted misinformation campaigns or personalized political advertising. [^19] See, for example, Collier et al. (2022) and Andrić and Kasirzadeh (2023). AI-infused surveillance and prediction technologies could create a self-reinforcing cycle: as governments increase surveillance, citizens resist these measures, leading to more intensive surveillance and control, ultimately weakening democratic institutions. [^20] See Manheim and Kaplan (2019), Crawford (2021), Madan and Ashok (2023), and Schaake (2024), among others.
The military subsystem, deeply intertwined with both economic and political subsystems, represents another critical node. AI’s integration into defense and intelligence operations is core to national security strategies, while simultaneously influencing international relations. Military applications of AI could create feedback loops that affect both technological development priorities and geopolitical power dynamics. [^21] See Scharre (2018), Morgan et al. (2020), and Zuboff (2019), among others.
The three subsystems do not operate in isolation but form a densely interconnected network where changes in one area compound through others, potentially in unexpected ways. Collectively, these subsystems, along with others, create the context within which AI x-catastrophes and their associated AI x-risks must be analyzed.
@@ 3 Systems analysis for AI x-risk
The next section applies this systems perspective to develop an illustrative case for the accumulative AI x-risk hypothesis: “The perfect storm MISTER” thought experiment. [^22] The term ”perfect storm” refers to a situation where a rare combination of circumstances drastically aggravates an event. It is typically used to describe scenarios in which a confluence of factors or events, which are individually manageable, come together to create an extraordinary and often catastrophic situation. This term entered the popular lexicon following the success of ”The Perfect Storm,” a 2000 American biographical disaster drama film directed by Wolfgang Petersen, adapted from Sebastian Junger’s 1997 non-fiction book. The story in both the book and film recounts a catastrophic weather event, where multiple meteorological elements converged to create a fierce and fatal storm. In broader applications, particularly in discussing systems or societal issues, ”perfect storm” characterizes situations where diverse negative factors or risks coalesce and interact. The interplay of these elements produces a compounded impact, far surpassing the severity one would expect from the sum of the individual parts. This convergence leads to a critically severe situation, often characterized by its heightened difficulty in management and resolution. The ”perfect storm” metaphor hence captures the essence of scenarios where the convergence of various challenges or risks creates a crisis of extraordinary magnitude. Each letter in MISTER represents a member of a subset of the social risks (Manipulation, Insecurity threats, Surveillance and erosion of Trust, Economic destabilization, and Rights infringement) that were introduced in Section 2.3.232323While not all social or ethical risks have existential implications, the Perfect Storm MISTER scenario focuses on those with potential for significant systemic impact. This will be followed by a comparison of decisive and accumulative AI x-risk pathways from a systems analysis perspective.